OpenStudy (anonymous):

Find f(-11/5) if f(n) = 5n + 6.

OpenStudy (anonymous):

Make n= -11/5 in the equation 5n+6 and solve.

OpenStudy (anonymous):

\[\LARGE \color{blue}{f(n)=5n+6}\] \[\LARGE \color{blue}{f(-\frac{11}{5})=5(-\frac{11}{5})+6 = -11 +6 = -5}\]